# Collision Energy Dependence of Particle Ratios and Freeze-out Parameters in Ultra Relativistic Nucleus Nucleus Collisions

### v1

@@ Abstract
This work investigates how the theoretically calculated particle ratios and freeze-out parameters for a multi-component hot and dense non-ideal hadron resonance gas system, formed during the pre freeze-out stage in ultra-relativistic nucleus-nucleus collisions, are affected by van der Waals type interactions, incorporating both repulsive and attractive interactions among hadrons. The baryons (antibaryons) are treated as hard incompressible objects. Using this theoretical approach, we analyze experimental particle ratios involving like-mass as well as unlike-mass hadrons. We also try to understand the existence of horn-like structures seen in the ratios of strange particles to pion as a function of the collision energy. Using information about the dependence of experimental particle ratios on $\sqrt{s_{NN}}$ over a wide range of collision energies, we have extracted several model freeze-out parameters and examined their dependence on $\sqrt{s_{NN}}$ The correlation between the $\bar{p}/p$ and $K^{-}/K^{+}$ ratios is also examined.
@@ I INTRODUCTION
Understanding the quark-hadron phase transition diagram and the nuclear matter under extreme conditions also requires investigating the thermodynamics of the hot and dense matter created in ultra-relativistic nucleus-nucleus collisions (URNNC). The URNNC are a key tool for experimentally studying the high energy-density and high-temperature states of matter, such as the Quark-Gluon Plasma (QGP). These collisions provide insights into collective effects, strange hadron yields and the possibility of a potential phase transition from a hadron resonance gas (HRG) phase, where quarks and gluons are confined to a phase where they are deconfined or chirally symmetric phase. The URNNC programs at the Relativistic Heavy Ion Collider (RHIC) at BNL and the Large Hadron Collider (LHC) at CERN aim to study the properties of strongly interacting matter under extreme conditions. They also offer a unique opportunity to study the chemical equilibration of hadrons in the final state of collisions Muller1995 ; Adcox2005 ; Singh1992 ; Tannenbaum2006 ; Cleymans2011 .
@@
Understanding the behavior of matter a few microseconds after the Big Bang in the early universe and within the core of the neutron stars has gained significant attention. In URNNC the exploration of strongly interacting matter at extreme temperatures and densities is a focal point for both experimental and theoretical investigations. At the laboratory level experimental endeavors such as those conducted at the Large Hadron Collider (LHC) and the Relativistic Heavy Ion Collider (RHIC) play a crucial role in unraveling the mysteries of hot and dense hadronic matter. The study of URNNC has shown that statistical-thermal models can quite successfully reproduce the essential particle production features and their relative multiplicities in such collisions. The statistical hadronization model (SHM) also reproduces transverse momentum spectra of different particle species, making it a useful tool for analyzing soft hadron production, particularly strangeness production which has been proposed as a signature of QGP formation Metag1993 ; Cleymans1986O ; Letessier1993 ; Becattini1998 ; Munzinger1999 ; Yen1998 ; Parra2019S .
@@
Theoretical models, particularly the HRG model, have been instrumental in providing insights into the complex dynamics of strongly interacting particles. The Ideal Hadron Resonance Gas (IHRG) model, an initial formulation, treats mesons and baryons as point-like entities, assuming no interaction among them. However, deviations from an uncorrelated gas of hadrons are observed in lattice Quantum Chromodynamics (QCD) simulations Becattini2000 ; Munzinger2001 ; Florkowski2001 ; Becattini1995 ; Cleymans1999 ; Munzinger1994 . The ideal excluded volume type HRG models play a crucial role in understanding the quark-hadron phase transition and the particle production in URNNC Andronic2006 . In a Monte Carlo type approach with excluded-volume effect taken into account in the HRG, calculation of arbitrary moments of the event-by-event hadron yields have been done. However an analysis of different types of particle ratios is still lacking Bazavov2012 . It nevertheless highlights the systematic excess of experimental pion abundances compared to ideal gas results, which attempted by assigning the pions a small hard-core volume Borsanyi2014 . These can lead to further review the various thermal and statistical descriptions of particle production in URNNC, including the excluded-volume effect in the HRG model.
@@
For hot and dense hadron gas, several approaches have been proposed to formulate a precise equation of state (EOS), often using phenomenological models due to the lack of lattice QCD results at finite baryon density $(n_{B})$. Thermal models, assuming thermo-chemical equilibrium, are effective in characterizing HRG, with the IHRG model being a basic version that treats hadrons and resonances as non-interacting point-like particles at freeze-out Welke1990 . It has been established that the short-range repulsion must be included to accurately describe HRG behaviour, as the IHG model’s phase transition calculations reveal inconsistencies at high temperatures and baryon densities Cleymans1986O ; Cleymans1986Q . The hadronic hard-core repulsive interactions become significant in a dense HRG and hence in a macroscopic approach this can be taken into account by assigning hadrons a geometrical hard-core volume, resulting in a van der Waals excluded-volume effect that reduces particle densities in the system. Recent phenomenological models incorporating this effect Singh1991 ; Cleymans1987I ; Kouno1989 ; Rischke1991 ; Anchishkin1992 ; Mir2023 ; Uddin2012 ; Parra2019 ; Mir2024 ; Bashir2018I ; Parra2022M often lack a well-defined partition function or thermodynamic potential ($\Omega$), leading to thermodynamical inconsistencies in the calculations of particle densities, i.e. $n\neq\partial\Omega/\partial\mu$, where $\mu$ is the chemical potential of the particle and. It also breaks causality in dense HRG, making them less appealing despite their some empirical success.
@@
Some earlier studies have also highlighted the importance of studying particle ratios as a function of the centre of mass energy in the URNNC. Hanafy (2021) Hanafy2021 found that a statistically corrected HRG model can reasonably describe experimental results across different energies. Tawfik (2012) Tawfik2012P also emphasized the significance of these ratios, particularly in the context of dynamical fluctuations in URNNC. Earlier Bjorken (1974) Bjorken1974 had also underscored the relevance of these ratios in understanding particle production in energetic collisions Blume2011 ; Tiwari2011 ; Tawfik2017 ; Stiles2006 ; Mir2024 ; Andronic2005 ; Bashir2015 ; Chatterjee2015 ; Mishra2008 ; Tiwari2013 .
@@
The interesting horn-like structure in the kaon to pion ratio, as observed in URNNC, has been attributed to the release of a large number of colour degrees of freedom Nayak2005 ; Nayak2010 ; Nayak2011 . This phenomenon has been further explored in a hadronic kinetic model, which successfully reproduces the sharp peak in the $K^{+}/\pi^{+}$ ratio Tomik2006 . These findings collectively suggest that the horn in the kaon to pion ratio is a result of the complex interplay of colour degrees of freedom and collision processes. However, in the present study we have attempted to investigate in detail this characteristic horn structure in the kaon-to-pion ratio as well as other strange hadron to pion ratios in URNNC. Within the framework of the model used in the present study we have found strong evidence of strangeness imbalance factor, $\xi_{s}$. This helps us in understanding the production of strange particles relative to pions in the system produced in URNNC Bir1982 ; Grebieszkow2009 ; Bhattacharyya2017 ; Tawfik2021H ; Cleymans2017T ; Oeschler2017 ; Cleymans2004T .
@@
The present work is organized as follows. In Section II we have provided some basic idea of the model formulation. In Section III the results of our calculations and their comparison with the experimental data is presented along with the discussion. Finally in Section IV we have summarized our results.
@@ II STATISTICAL THERMAL MODEL
The symbol $n$ represents the finite-size particle number density. The first term includes the correction factor due to hard-core repulsion by assigning a hard-core radius to the particles (baryons and antibaryons in our case). Considering particles as rigid spheres, we have $b=16\pi r^{3}/3$, where $r$ is the particle’s hard-core radius. The second term accounts for the attractive interactions between particles through the parameter $a$. For $a=0$, the equation reduces to the EV-HRG equation, containing only repulsive interactions, while for both $a=0$ and $b=0$ it reduces to the ideal HRG case Greiner1965 ; Tiwari2011 ; Bugaev2000 ; Sahoo2023 ; Yen1997 .
In a statistical approach, it is necessary to adopt the grand canonical ensemble (GCE) formulation for the system Samanta2017C ; Vovchenko2018Q ; Vovchenko2017M ; Pradhan2022 ; Vovchenko2015P ; Sarkar2018 ; Vovchenko2015S ; Vovchenko2015V . The preference for the Grand Canonical Ensemble (GCE) over the Canonical Ensemble (CE) stems from its technical simplicity in describing the properties of the systems due to its non-conserved (fluctuating) particle numbers and energy. Hence, the GCE is most suited for the HRG system formed in the URNNC essentially due to the particle production/annihilation reactions taking place inside it. The CE formulation thus lacks a comprehensive thermodynamic description for such systems due to its dependence on parameters like volume ($V$), temperature ($T$), energy ($E$), and particle number ($N$). In the GCE formulation, we can calculate mean particle number density ($n$), energy density ($\varepsilon$), entropy density ($s$), etc., which are essential for understanding the system’s behaviour. Here we will apply the following statistical approach, which was earlier performed for the EV model, i.e., for $a=0$ in Eq. (2), Rischke1991 ; Gorenstein1981 ; Parra2019 ; Parra2023 ; Parra2019T ; Din2024 .
II.1 VDW EQUATION OF STATE THROUGH GRAND CANONICAL ENSEMBLE:
@@
The summation over the index $i$ is for all species having hard-core repulsive interaction, including the $j$th specie. Further details can be found in Parra2019 . The above results can be effectively utilized to calculate relative particle yields in the system.
@@
Where $\Gamma_{k\rightarrow j}$ is the branching ratio for the decay of the $k^{th}$ resonance into the $j^{th}$ particle state. The superscript $prim$ stands for the primary number density without decay contribution. The summation is over all the heavier resonances $k$ that decay to the $j$th hadron. Here we have considered the decay contributions to the final state hadron multiplicities from single weak decays of the heavy hadronic resonances as well as those double decays where a weak decay is followed by a strong decay.
The parameters $a$ and $b$ of the model can be adjusted to reproduce the ground state properties of cold nuclear matter (i.e., $T\approx 0\,\text{MeV}$) Vovchenko2015V ; Bethe1971 ; Cleymans1986Q ; Wang2022 ; Greiner1965 ; Bugaev2000 ; Zhang1992 by using the binding energy $E_{B}\approx 16\,\text{MeV}$ and the nucleon number density $n=n_{0}\approx 0.16\,\text{fm}^{-3}$. Following this, one finds $a=329\,\text{MeV}\,\text{fm}^{3}$ and $b=3.42\,\text{fm}^{3}$, which gives the nucleonic hard-core radius $r=0.59\,\text{fm}$, a very reasonable value Vovchenko2016V .
III RESULTS AND DISCUSSION
In this section, we present the outcome of our analysis of particle production, including the strangeness imbalance effect observed in URNNC over a wide range of collision energy, using the above discussed VDW-HRG model. For the $i^{\text{th}}$ hadron having baryon number $B_{i}$, strangeness $S_{i}$, and electric charge $Q_{i}$, we associate baryon chemical potential $\mu_{B}$, strange chemical potential $\mu_{S}$, and electric chemical potential $\mu_{Q}$, respectively. These are treated as model parameters which control the net baryon, strangeness, and electric charge content of the system. The overall chemical potential of the $i^{\text{th}}$ hadron is thus defined as:
@@ III RESULTS AND DISCUSSION
As protons (antiprotons) are the lightest and most abundantly produced baryons (antibaryons) in the system hence we have first attempted to determine the parameters in Eq. 35 and 36 from the best fit to the experimental data on the $\bar{p}/p$ ratio as a function of $\sqrt{s_{NN}}$ in URNNC . The freeze-out temperature has been found to be almost universal at RHIC and LHC, thereby suggesting that chemical freeze-out might occur soon after hadronization Bashir2015 . These studies highlight the complex interplay of factors in determining freeze-out parameters, hence requiring detailed analysis. To ascertain the values of the parameters ($c,d,e,f,g$), we use the best fit criteria, i.e., by obtaining the minimum value of the chi-squared per degree of freedom ($\chi^{2}/\text{dof}$), to the experimentally measured $\bar{p}/p$ ratio distribution over a wide range of collision energies, $\sqrt{s_{NN}}$. The chi-square ($\chi^{2}/\text{dof}$) is computed as:
@@
where $R_{i}^{\text{the}}$ and $R_{i}^{\text{exp}}$ represent theoretical and experimental values, respectively, with $\sigma_{i}$ as the experimental uncertainties in the measured quantities at corresponding $\sqrt{s_{NN}}$, with $N$ as the number of degrees of freedom in the data set. The values of the constants turn out to be:
$c=1402\,\text{MeV},\quad d=0.26\,\text{GeV}^{-1},\quad e=172\,\text{MeV},$
$f=0.34\,\text{MeV}^{-1},\quad\text{and}\quad g=0.015\,\text{MeV}^{-3}$
resulting in a satisfactory fit with a value of $\chi^{2}/\text{dof}=1.67$ with N = 15. The $\sqrt{s_{NN}}$ dependent theoretical curve for the $\bar{p}/p$ obtained using the VDW-HRG model calculations is shown in Fig. 1 . Experimental data has been taken from various available sources Adamczyk2017B ; Ahle1999P ; Adams2002S ; Aggarwal2011S ; Aamodt2010M .
Figure 1: $\bar{p}/p$ dependence on $\sqrt{s_{NN}}$
III.1 LIKE MASS ANTIHYPERON TO HYPERON RATIOS

### v2

@@ Abstract
This work investigates the thermo-chemical freeze-out condition of the multi-component hot and dense hadron resonance gas (HRG) formed in the ultra-relativistic nucleus-nucleus collisions (URNNC). The van der Waals (VDW) type model used in the present analysis incorporates the repulsive as well as attractive interactions among the hadrons. The baryons (antibaryons) are treated as incompressible objects. Using this theoretical approach the values of the model freeze-out parameters of the system are extracted over a wide range of collision energy by analyzing experimental data on like-mass antibaryon to baryon ratios. The same set of parameters is found to explain the energy dependence of several other particle ratios quite satisfactorily. We find that the horn-like structures seen in the ratios of strange particles to pions as a function of the collision energy cannot be explained by the VDW-HRG model alone without considering the strangeness imbalance effect in the system. We have compared our freeze-out line with those obtained earlier. The correlation between the $\bar{p}/p$ and $K^{-}/K^{+}$ ratios is also examined.
@@ I INTRODUCTION
Understanding the quark-hadron phase transition diagram and the nuclear matter under extreme conditions requires investigating the thermodynamics of the hot and dense matter created in ultra-relativistic nucleus-nucleus collisions (URNNC). Exploring the behavior of matter a few microseconds after the Big Bang in the early universe and within the core of the neutron stars has gained significant attention.
@@
The URNNC are a key tool for experimentally studying the high energy-density and high-temperature states of matter, such as the Quark-Gluon Plasma (QGP) in the laboratory. These collisions provide insights into collective effects, strange hadron yields and the possibility of a potential phase transition from a hadron resonance gas (HRG) phase, where quarks and gluons are confined, to a phase where they are deconfined or in a chirally symmetric phase. The URNNC programs at the Relativistic Heavy Ion Collider (RHIC) at BNL and the Large Hadron Collider (LHC) at CERN aim to study the properties of strongly interacting matter under extreme conditions. They also offer a unique opportunity to study the chemical equilibration of hadrons in the final state of collisions Muller1995 ; Adcox2005 ; Singh1992 ; Tannenbaum2006 ; Cleymans2011 .
@@
In URNNC the exploration of strongly interacting matter at extreme temperatures and densities is a focal point for both experimental and theoretical investigations. Such studies have shown that statistical-thermal models can quite successfully reproduce the essential particle production features including their spectra in such collisions. The statistical hadronization model (SHM) reproduces transverse momentum spectra of different particle species, making it a useful tool for analyzing soft hadron production, particularly strangeness production which has been proposed as a signature of QGP formation Metag1993 ; Cleymans1986O ; Letessier1993 ; Becattini1998 ; Munzinger1999 ; Yen1998 ; Parra2019S .
@@
Theoretical models, particularly the HRG model, have been instrumental in providing insights into the complex dynamics of strongly interacting particles. The Ideal Hadron Resonance Gas (IHRG) model, an initial formulation, treats mesons and baryons as point-like entities, assuming no interaction among them. However, deviations from an uncorrelated gas of hadrons are observed in lattice Quantum Chromodynamics (QCD) simulations Becattini2000 ; Munzinger2001 ; Florkowski2001 ; Becattini1995 ; Cleymans1999 ; Munzinger1994 . The ideal excluded volume type HRG models play a crucial role in understanding the quark-hadron phase transition and the particle production in URNNC Andronic2006 . In a Monte Carlo type approach with excluded-volume effect taken into account in the HRG, calculation of arbitrary moments of the event-by-event hadron yields have been done. However explaining the different types of particle ratios simultaneously over a wide range of energy is still lacking Bazavov2012 . It has also highlighted the systematic excess of experimental pion abundances when compared to the ideal hadron resonance gas results. This has also been attempted by assigning the pions a finite size Borsanyi2014 . These can lead to further review the various thermal and statistical descriptions of particle production in URNNC, including the excluded-volume effect in the HRG model.
@@
For hot and dense hadron gas, several approaches have been proposed to formulate a precise equation of state (EOS), often using phenomenological models due to the lack of lattice QCD results at finite baryon density $(n_{B})$. Thermal models, assuming thermo-chemical equilibrium, are effective in characterizing HRG, with the IHRG model being a basic version that treats hadrons and resonances as non-interacting point-like particles at freeze-out Welke1990 . It has been established that the short-range repulsion must be included to accurately describe HRG behaviour, as the IHRG model’s phase transition calculations reveal inconsistencies at high temperatures and baryon densities Cleymans1986O ; Cleymans1986Q . The hadronic hard-core repulsive interactions become significant in a dense HRG and hence in a macroscopic approach this can be taken into account by assigning hadrons a geometrical hard-core volume, resulting in a van der Waals excluded-volume type effect that reduces particle densities in the system. Phenomenological models incorporating this effect Singh1991 ; Cleymans1987I ; Kouno1989 ; Rischke1991 ; Anchishkin1992 ; Mir2023 ; Uddin2012 ; Parra2019 ; Mir2024 ; Bashir2018I ; Parra2022M often lack a well-defined partition function or thermodynamic potential ($\Omega$), leading to thermodynamical inconsistencies in the calculations of particle densities, i.e. $n\neq\partial\Omega/\partial\mu$, where $\mu$ is the chemical potential of the particle. It also breaks causality in dense HRG, making them less appealing despite their some empirical success.
@@
Some earlier studies have also highlighted the importance of studying particle ratios as a function of the centre of mass energy in the URNNC. Hanafy Hanafy2021 found that a statistically corrected HRG model can reasonably describe experimental results across different energies. Tawfik Tawfik2012P also emphasized the significance of these ratios, particularly in the context of dynamical fluctuations in URNNC. Earlier Bjorken Bjorken1974 had also underscored the relevance of these ratios in understanding particle production in energetic collisions Blume2011 ; Tiwari2011 ; Tawfik2017 ; Stiles2006 ; Mir2024 ; Andronic2005 ; Bashir2015 ; Chatterjee2015 ; Mishra2008 ; Tiwari2013 .
@@
The interesting horn-like structure in the kaon to pion ratio, as observed in URNNC, has been attributed to the release of a large number of colour degrees of freedom Nayak2005 ; Nayak2010 ; Nayak2011 ; Andronic2009Horn . This phenomenon has been further explored in a hadronic kinetic model, which successfully reproduces the sharp peak in the $K^{+}/\pi^{+}$ ratio Tomik2006 . These findings collectively suggest that the horn in the kaon to pion ratio is a result of the complex interplay of colour degrees of freedom and collision processes. However, in the present study we have attempted to investigate in detail this characteristic horn structure in the kaon-to-pion ratio as well as other strange hadron to pion ratios in URNNC. Within the framework of the model used in the present study we have found strong evidence of strangeness imbalance factor, $\xi_{s}$. This helps us in understanding the production of strange particles relative to pions in the system produced in URNNC Bir1982 ; Grebieszkow2009 ; Bhattacharyya2017 ; Tawfik2021H ; Cleymans2017T ; Oeschler2017 ; Cleymans2004T .
@@
The present work is organized as follows. In Section II and  III we have provided some basic idea of the model formulation. In Section IV the results of our calculations and their comparison with the experimental data are presented along with the discussion. Finally in Section V we have summarized our results.
@@ II STATISTICAL THERMAL MODEL
The symbol $n$ represents the finite-size particle number density. The first term includes the correction factor due to hard-core repulsion by assigning a hard-core radius to the particles (baryons and antibaryons in our case). Considering particles as rigid non-deformable spheres, we have $b=16\pi r^{3}/3$, where $r$ is the particle’s hard-core radius. The second term accounts for the attractive interactions between particles through the parameter $a$. For $a=0$, the equation reduces to the EV-HRG equation, containing only repulsive interactions, while for both $a=0$ and $b=0$ it reduces to the ideal HRG case Greiner1965 ; Tiwari2011 ; Bugaev2000 ; Sahoo2023 ; Yen1997 .
In a statistical approach, it is necessary to adopt the grand canonical ensemble (GCE) formulation to study the systems formed in the URNNC Samanta2017C ; Vovchenko2018Q ; Vovchenko2017M ; Pradhan2022 ; Vovchenko2015P ; Sarkar2018 ; Vovchenko2015S ; Vovchenko2015V . The preference for the Grand Canonical Ensemble (GCE) over the Canonical Ensemble (CE) stems from its technical simplicity in describing the properties of the systems due to its non-conserved (fluctuating) particle numbers in the system which is essentially due to the particle production/annihilation reactions taking place inside it. The GCE approach is used to calculate the mean particle densities in the HRG system. The CE formulation on the other hand depends on parameters like volume ($V$), temperature ($T$), energy ($E$) and the particle number ($N$) of the system. In the GCE formulation, we can calculate mean particle number density ($n$), energy density ($\varepsilon$), entropy density ($s$), etc., which are essential for understanding the system’s behaviour. Here we will apply the following statistical approach, which was earlier used for the EV model, i.e., for $a=0$ in Eq. (2), Rischke1991 ; Gorenstein1981 ; Parra2019 ; Parra2023 ; Parra2019T ; Din2024 .
III VDW EQUATION OF STATE THROUGH GRAND CANONICAL ENSEMBLE:
@@
The summation over the index $i$ is for all species having hard-core repulsive interaction, including the $j^{th}$ specie. Further details can be found in Parra2022M ; Vovchenko2016V ; Vovchenko2015P . The above results can be effectively utilized to calculate relative particle yields in the system.
@@
Where $\Gamma_{k\rightarrow j}$ is the branching ratio for the decay of the $k^{th}$ resonance into the $j^{th}$ particle state. The superscript $prim$ stands for the primary number density without decay contribution. The summation is over all the heavier resonances $k$ that decay to the $j^{th}$ hadron. Here we have considered the decay contributions to the final state hadron multiplicities from single weak decays of the heavy hadronic resonances as well as those double decays where a weak decay is followed by a strong decay.
In the present analysis we have treated b as a free parameter which is fitted to observed particle ratios, yielding values in close agreement with estimates based on lattice QCD data. Tiwari2011 ; Samanta2017C ; Vovchenko2017E ; Albright2014M ; Mishra2007Effect ; Sarkar2018 . The value of the parameter $a=329\,\text{MeV}\ \text{fm}^{3}$ of the model is chosen from previous works Vovchenko2015V ; Bethe1971 ; Cleymans1986Q ; Wang2022 ; Greiner1965 ; Bugaev2000 ; Zhang1992
IV RESULTS AND DISCUSSION
In this section, we present the outcome of our analysis of particle production, including the strangeness imbalance effect observed in URNNC over a wide range of collision energy, using the above discussed VDW-HRG model.
For the $i^{\text{th}}$ hadron having baryon number $B_{i}$, strangeness $S_{i}$, and electric charge $Q_{i}$, we associate baryon chemical potential $\mu_{B}$, strange chemical potential $\mu_{S}$, and electric chemical potential $\mu_{Q}$, respectively. These are treated as model parameters which control the net baryon, strangeness, and electric charge content of the system. The overall chemical potential of the $i^{\text{th}}$ hadron is thus defined as:
@@ IV RESULTS AND DISCUSSION
To ascertain the values of the parameters ($c,d,e,f,g$), we use the best fit criteria, i.e., by obtaining the minimum value of the chi-squared per degree of freedom ($\chi^{2}/\text{{dof}}$), to the experimentally measured antibaryon to baryon ratios over a wide range of collision energy, $\sqrt{s_{NN}}$. The chi-square ($\chi^{2}/\text{{dof}}$) is computed as:
@@
where $R_{i}^{\text{the}}$ and $R_{i}^{\text{exp}}$ represent theoretical and experimental values, respectively, with $\sigma_{i}$ as the experimental uncertainties in the measured quantities at corresponding $\sqrt{s_{NN}}$, with $N$ as the number of degrees of freedom in the data set. The values of the model parameters are extracted from the best fit method and are found to be :
$c=1398\pm 29\,\text{MeV},\quad d=0.28\pm 0.008\,\text{GeV}^{-1},$
$e=164\pm 2.5\,\text{MeV},\quad f=0.14\pm 0.021\,\text{MeV}^{-1},$
$g=0.015\pm 0.04\,\text{MeV}^{-3}$
The values of the above parameters along with that of the baryonic hard-core radius, r = $0.73\pm 0.02\,\text{fm}$, provide (nearly) simultaneous minimum $\chi^{2}/\text{{dof}}$ fit to all four antibaryon to baryon ratios (i.e., $\bar{p}/p$, Λ¯/$\bar{\Lambda}/\Lambda$, $\bar{\Xi}^{-}/\Xi^{-}$ and Ω¯/$\bar{\Omega}/\Omega$). This estimate of the baryonic hard-core radius is in close agreement with the earlier estimate based on the description of the lattice QCD data of different thermodynamics quantities Samanta2017C .
IV.1 LIKE MASS ANTIBARYON TO BARYON RATIOS